OpenStudy (anonymous):

-5y - 9=-(y - 1) solve please real quick need answer :)

OpenStudy (anonymous):

Solve for y: -9-5 y = 1-y Move terms with y to the left hand side. Add y to both sides: -(5 y)+y-9 = (-y+y)+1 Look for two terms that sum to zero. -y+y = 0: -(5 y)+y-9 = 1 Move everything to the left hand side. Subtract 1 from both sides of y-5 y-9 = 1: -(5 y)+y-9-1 = 1-1 Look for two terms that sum to zero. 1-1 = 0: -(5 y)+y-9-1 = 0 Add the numbers in -(5 y)+y-9-1 together. Add like terms. -9-1 = -10: -5 y+y-10 = 0 Factor the polynomial, -5 y+y-10. Factor -2 from the polynomial -4 y-10: -2 (5+2 y) = 0 Divide both sides by a constant to simplify the equation. Divide both sides of -2 (2 y+5) = 0 by -2: (-2 (2 y+5))/(-2) = 0/(-2) Any nonzero number divided by itself is one. (-2)/(-2) = 1: 2 y+5 = 0/(-2) Any number times zero is zero. 0/(-2) = 0: 2 y+5 = 0 Isolate terms with y to the left hand side. Subtract 5 from both sides: 2 y+(5-5) = -5 Look for two terms that sum to zero. 5-5 = 0: 2 y = -5 Solve for y. Divide both sides of 2 y = -5 by 2: (2 y)/(2) = -5/2 Any nonzero number divided by itself is one. 2/2 = 1: Answer: | | y = -5/2

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Click best response if this helped!